OpenStudy (anonymous):

Simplify the following expression. a + 2c - 5b - c + a - b

OpenStudy (mathstudent55):

Combine like terms. Like terms have the same variable part (the letter part). a and a are like terms 2c and -c are like terms -5b and -b are like terms A term written with no number, is as if the number 1 is there: a = 1a -a = -1a To combine like terms, you add the number parts together and write the variable part afterwards. For example: 2a + 3a = 5a 2b - b = 2b - 1b = 1b = b

OpenStudy (anonymous):

thanks this helped